# MVP-ORAM: a Wait-free Concurrent ORAM for Confidential BFT Storage

### v1

@@ X Conclusions and Future Work
This paper presented MVP-ORAM, the first Byzantine fault-tolerant ORAM protocol. It enables fail-prone concurrent clients to access a shared data store without revealing any information about the accessed data or their accesses. We show that in asynchronous networks, satisfying wait freedom fundamentally compromises collision freedom, affecting the security guarantees of our construction. To account for that, we propose a weaker security definition for asynchronous wait-free ORAMs, which might be secure enough for typical storage applications, with skewed block accesses. We devise MVP-ORAM as a deterministic wait-free ORAM service and integrate it into a confidential BFT data store, which shows promising performance results. Additionally, we introduce a stronger variant of MVP-ORAM that ensures perfect access-pattern secrecy with additional assumptions.
This paper opens many avenues for future work. For example, it seems impossible to implement a perfect wait-free ORAM in asynchronous systems, but this remains to be proved. Further, more efficient variants of MVP-ORAM can be devised for implementing perfect obliviousness. Finally, state-of-the-art information dispersal [OptimalAVID] can potentially be used to improve the bandwidth requirements and concrete performance of replicated/BFT ORAM.
@@ Acknowledgment
We thank the NDSS 2026 anonymous reviewers for their constructive comments, which helped improve the paper. We also thank Cristiano Santos for his initial work on BFT ORAM services, which sparked the results presented in this paper. This work was supported by FCT through the Ph.D. scholarship, ref. 2020.04412.BD, (https://doi.org/10.54499/2020.04412.BD), the SMaRtChain project, ref. 2022.08431.PTDC (https://doi.org/10.54499/2022.08431.PTDC), and the LASIGE Research Unit, ref. UID/408/2025.
@@ Appendix A MVP-ORAM Auxiliary Functions
Algorithm 3 specifies the auxiliary functions used by clients to update their position map ($\mathsf{consolidatePathMaps}$), merge multiple versions of path and stashes ($\mathsf{mergePathStashes}$), and to create a new version of path and stash ($\mathsf{populatePath}$).
@@ B-A Preliminary Definitions
ORAM access operations. MVP-ORAM requires every operation to be executed in the same order in every server. When replicated, this order is defined by the underlying consensus (or total order broadcast) protocol implemented by the BFT SMR. Due to this, every operation invoked in the “logical” MVP-ORAM server can be totally ordered based on their execution order.
@@
Operation histories. The interaction between a set of clients and our replicated state machine is modeled by a sequence of operation invocation and reply events called history [herlihy_1990]. A history $\mathcal{H}=(H,<_{H})$ comprises the set of events $H$ and a total order relation among these events $_{H}$. The order of events in a history $\mathcal{H}$ defines a partial order of operations $\rightarrow_{H}$. We say an operation $o_{1}$ precedes (resp. succeeds) an operation $o_{2}$, denoted by $o_{1}\rightarrow_{H}o_{2}$ (resp. $o_{2}\rightarrow_{H}o_{1}$), if $rep(o_{1})<_{H}inv(o_{2})$ (resp. $rep(o_{2})<_{H}inv(o_{1})$). If an operation neither precedes nor succeeds another, we say the two operations are concurrent.
@@ B-C Safety: Memory Linearizability
If $w=\mathsf{access}(*,write,b,x)$ is the closest preceding write of read $r=\mathsf{access}(*,read,$ $b,\bot)$, then $r$ returns $x$.
@@
For each memory position $b$, MVP-ORAM’s read ($\mathsf{access}(*,read,b,\bot)$) and write ($\mathsf{access}(*,write,$ b,)b,*)) operations satisfy linearizability.
@@ Appendix C Obliviousness Proofs
The access pattern of MVP-ORAM depends on the number of concurrent clients and the distribution of the accessed blocks, and it can be described by three scenarios: (1) a single client accesses the ORAM once per timestep, (2) multiple clients access different blocks per timestep, and (3) multiple clients access the same block in the same timestep. In practice, an access pattern will be a mixture of those scenarios, i.e., in a timestep, a single client accesses a block, while in another timestep, $c_{d}$ clients access distinct blocks and $c-c_{d}$ clients access the same block. However, we analyze those scenarios separately to show the best- and worst-case security of MVP-ORAM. The first two scenarios represent the best-case scenarios as clients access different blocks in each timestep, which generates random leaves. However, the generated leaves in the third scenario are correlated, which is why that scenario represents the worst case.
@@
  • If $b_{i}\neq b_{j}$. Note that the only scenario in which $x_{i}$ could be connected to $x_{j}$ in this case is as follows: during the $i$th access, $b_{i}$ is moved to the stash, $b_{j}$ is already in the stash, and $b_{j}$ is subsequently moved to the path $\mathcal{P}(x_{i})$. Although this connection exists, due to F3, $b_{j}$ is placed close to the root, allowing $x_{j}$ to be almost any leaf. Consequently, $x_{i}$ remains independent of $x_{j}$.
@@ 29 $\mathcal{S}^{*}\leftarrow(\mathcal{S}^{c}\setminus\mathcal{S})\cup\{S\}$
The Strong MVP-ORAM protocol makes access pattern oblivious even in the worst-case scenario when ORAM accesses follow a uniform distribution by executing $\sigma\geq 0$ dummy accesses for each real access. The proposed approach creates a scheduler of accesses containing $\sigma$ dummy accesses and a single real access, such that in each timestep, no two clients will access the same block if $c\leq\sigma+1$. To construct this scheduler, each client specifies the address it will access when invoking $\mathit{getPM}$, and they utilize this information to create the scheduler. Note that the address is encrypted; however, for simplicity, we omit the encryption details from the algorithm specification. The modified client and server specifications are presented in Algorithm 4 and Algorithm 5, respectively.
The implementation of this strategy required changing a few lines in the MPV-ORAM protocol. The first change is the addition of the list $\mathcal{A}$ in the server. For each client $c_{i}$, the server stores in $\mathcal{A}$ the encrypted address that $c_{i}$ will access and a counter that tracks the number of accesses executed by $c_{i}$ (A5, L7-10). The address is used by $c_{i}$ to determine the timestep at which it will perform the real access, while the counter is used by the server to remove $c_{i}$’s access context on the $(\sigma+1)^{\mathit{th}}$ access’s eviction (A5, L18-20).
@@
Using $\tau_{\mathit{real}}$, $c_{i}$ verifies twice whether the current timestep $\tau$ matches $\tau_{\mathit{real}}$ during the access. If the check returns true, $c_{i}$ executes the real access by selecting a random path that passes through the slot where the block is stored (A4, L9-10) and applying the operation of $\mathit{read}$ or $\mathit{write}$ on the block (A4, L15-20). Otherwise, it selects a random path (A4, L12) to execute dummy access. Note that if the number of concurrent clients $c$ accessing the same address is more than $\sigma+1$, then $c-\sigma$ clients will have to execute real access in the same last timestamp. Hence, if $c\leq\sigma+1$, then this strategy guarantees obliviousness even when the worst-case access pattern, i.e., all clients are concurrently accessing the same address, is not skewed. Next, we prove this result.
@@ E-A Obliviousness of Strong MVP-ORAM
Given $c,\sigma,N\in\mathbb{N}$, $\alpha\in\mathbb{R}$, and $D\in\mathcal{U}$, the statistical distance between a random sequence of size $c$ and the access pattern generated by Strong MVP-ORAM under the dummy access strategy is bounded by $\mu(N,c,\mathcal{D}(\alpha),\sigma)$.
@@
  • If $c\leq\sigma+1$. Since clients can detect whether $\mathit{addr}$ is being accessed by others during timestep $t$, a client that is the $k$th to access $\mathit{addr}$, where $k\leq c$, performs k1k-1 dummy accesses before issuing its actual access. Indeed, the mitigation strategy ensures that accesses to $\mathit{addr}$ are performed so that no two clients access $\mathit{addr}$ in the same timestamp. As each block is moved to the stash at the end of every access (if it is not already there), each client selects a random, independent path when accessing the block. As a result, the access pattern generated by Strong MVP-ORAM is indistinguishable from a random sequence of size $c$.
@@
Given $c,\sigma\in\mathbb{N}$, let $\overrightarrow{b}_{e}=\{b_{i}\}_{i\in\{1,\dots,c\}}$ denote a set of $c$ concurrent queries in timestep $e$ and $\overrightarrow{y}=(\overrightarrow{b}_{1},\overrightarrow{b}_{2},\dots)$ denote a set of queries from multiple clients in each timestep. MVP-ORAM holds the Obliviousness property of a Wait-Free ORAM, as per Definition 1, when clients perform $\sigma$ dummy accesses.
@@
Let $A(\overrightarrow{y})$ be an access pattern generated by Strong MVP-ORAM while executing query $\overrightarrow{y}$. According to Definition 1, we must show that there is a function $\mu$ that bounds the statistical distance between any two access patterns of the same length. To establish such a bound, we compare $A(\overrightarrow{y})$ with a random access pattern. From Theorem 7, $A(\overrightarrow{y})$ is indistinguishable from a random access pattern of the same size when a single client accesses the ORAM in each timestep or multiple clients access different blocks. From Theorem 12, in the worst case, the statistical distance between $A(\overrightarrow{b}_{e})$ and a random access pattern of the same size is bounded by $\mu(N,c,\alpha,\sigma)$. Since after each concurrent access ORAM changes, i.e., locations of blocks accessed by clients change, the statistical distance between $A(\overrightarrow{y})$ and a random access sequence is still bounded by $\mu(N,c,\alpha,\sigma)$, proving the lemma. ∎
To conclude our obliviousness analysis, we show in which conditions MVP-ORAM can provide the same security as a traditional parallel ORAM with collision-freedom [boyle2015].
@@
When there are at most $\sigma+1$ clients accessing ORAM, they retrieve blocks in $\sigma+1$ timesteps. When they are retrieving different blocks, they will access different random paths in the first timestep and continue accessing random paths in the remaining $\sigma$ timesteps. When they retrieve the same block, each client will perform the real access in different timesteps and dummy accesses in the other ones. In both cases, clients generate a sequence of leaves by selecting any of the $2^{L}$ leaves in each access. Thus, the probability of all clients accessing the same leaf is $2^{-L\cdot c\cdot(\sigma+1)}$, which is the same probability of clients selecting random leaves and is negligible in $N$. ∎
@@ Appendix F Artifact Appendix
The artifact includes the implementation of the MVP-ORAM protocol and QuORAM [^4] https://github.com/SeifIbrahim/QuORAM/ for performance assessment and comparison, respectively. Experiments were conducted on multiple AWS servers, but a scaled-down version for single-machine execution is provided for artifact evaluation. Fig. 5 and Fig. 6 were generated through numerical simulations, which we do not include in the artifact evaluation.
@@ F-A1 How to access
The artifact is available on Zenodo [^5] https://doi.org/10.5281/zenodo.17842154, which includes the MVP-ORAM and QuORAM implementations adapted for easier benchmarking.
@@ F-A3 Software dependencies
Execution of this artifact requires unzip, gnuplot, and OpenJDK 11. It can be executed on either Linux or the Linux subsystem on Windows [^6] https://learn.microsoft.com/windows/wsl.
@@ F-D Major Claims
  • (C1): The stash size stabilizes after some number of accesses, and the maximum stash size increases as we decrease $\alpha$ and increase $c$. This is proven by experiments (E1), whose results are shown in Fig. 7.
@@
  • (C4): MVP-ORAM outperforms QuORAM both in throughput and latency with $50$ clients. This is proven by experiment (E3), with results reported in Table III.
@@ F-E1 Experiment (E1)
[Execution] Follow the steps described in the experimental workflow to execute the experiment twice, setting the parameters defined above.
@@ F-E2 Experiment (E2)
[Execution] Follow the steps described in the experimental workflow to execute the experiment using the above values.

### v2

@@ X Conclusions and Future Work
This paper presented MVP-ORAM, the first Byzantine fault-tolerant ORAM protocol. It enables fail-prone concurrent clients to access a shared data store without revealing any information about the accessed data or their access patterns. We show that in asynchronous networks, satisfying wait-freedom fundamentally compromises collision-freedom, affecting the security guarantees of our construction. To account for this, we propose a weaker security definition for asynchronous wait-free ORAMs, which may be secure enough for typical storage applications with skewed block accesses. We devise MVP-ORAM as a deterministic wait-free ORAM service and integrate it into a confidential BFT data store, which shows promising performance results. Additionally, we introduce a stronger variant of MVP-ORAM that ensures perfect access-pattern secrecy with additional assumptions.
This paper opens many avenues for future work. For example, it seems impossible to implement a perfect wait-free ORAM in asynchronous systems, but this remains to be proved. Furthermore, more efficient variants of MVP-ORAM can be devised to implement perfect obliviousness. Finally, state-of-the-art information dispersal [OptimalAVID] can potentially be used to improve the bandwidth requirements and concrete performance of replicated/BFT ORAM.
@@ Acknowledgment
We thank the anonymous reviewers for their constructive comments, which helped improve the paper. We also thank Cristiano Santos for his initial work on BFT ORAM services, which sparked the results presented in this paper. This work was supported by FCT through the Ph.D. scholarship, ref. 2020.04412.BD, the SMaRtChain and APOSTLE projects, ref. 2022.08431.PTDC and 2023.12254.PEX, respectively, and the LASIGE Research Unit, ref. UID/00408/2025.
@@ Appendix A MVP-ORAM Auxiliary Functions
Algorithm 3 specifies the auxiliary functions used by clients to update their position map ($\mathsf{consolidatePathMaps}$), merge multiple versions of path and stashes ($\mathsf{mergePathStashes}$), and to create a new version of path and stash ($\mathsf{populatePath}$), as used in Algorithm 1V).
@@ B-A Preliminary Definitions
ORAM access operations. MVP-ORAM requires every operation to be executed in the same order on every server. When replicated, this order is defined by the underlying consensus (or total order broadcast) protocol implemented by the BFT SMR. Due to this, every operation invoked in the “logical” MVP-ORAM server can be totally ordered based on its execution order.
@@
Operation histories. The interaction between a set of clients and our replicated state machine is modeled by a sequence of operation invocation and reply events called history [herlihy_1990]. For an operation $o$, we denote by $inv(o)$ its invocation event and by $rep(o)$ its corresponding reply event. A history $\mathcal{H}=(H,<_{H})$ comprises the set of events $H$ and a total order relation among these events $_{H}$. The order of events in a history $\mathcal{H}$ defines a partial order of operations $\rightarrow_{H}$. We say an operation $o_{1}$ precedes (resp. succeeds) an operation $o_{2}$, denoted by $o_{1}\rightarrow_{H}o_{2}$ (resp. $o_{2}\rightarrow_{H}o_{1}$), if $rep(o_{1})<_{H}inv(o_{2})$ (resp. $rep(o_{2})<_{H}inv(o_{1})$). If an operation neither precedes nor succeeds another, we say the two operations are concurrent.
@@ B-C Safety: Memory Linearizability
If $w=\mathsf{access}(\_,write,b,x)$ is the closest preceding write of read $r=\mathsf{access}(\_,read,$ $b,\bot)$, then $r$ returns $x$.
@@
For each memory position $b$, MVP-ORAM’s read ($\mathsf{access}(\_,read,b,\bot)$) and write ($\mathsf{access}(\_,write,b,\_)$) operations satisfy linearizability.
@@ Appendix C Obliviousness Proofs
The access pattern of MVP-ORAM depends on the number of concurrent clients and the distribution of the accessed blocks, and it can be described by three scenarios: (1) a single client accesses the ORAM once per timestep, (2) multiple clients access different blocks per timestep, and (3) multiple clients access the same block in the same timestep. In practice, an access pattern will be a mixture of these scenarios, i.e., in a timestep, a single client accesses a block, while in another timestep, $d$ clients access distinct blocks and cdc-d clients access the same block. However, we analyze those scenarios separately to show the best- and worst-case security of MVP-ORAM. The first two scenarios represent the best-case scenarios as clients access different blocks in each timestep, which generates random leaves. However, the generated leaves in the third scenario are correlated, which is why that scenario represents the worst-case.
@@
  • If $b_{i}\neq b_{j}$. Note that the only scenario in which $x_{i}$ could be connected to $x_{j}$ in this case is as follows: during the $i$th access, $b_{i}$ is moved to the stash, $b_{j}$ is already in the stash, and $b_{j}$ is subsequently moved to the path $\mathcal{P}_{x_{i}}$. Although this connection exists, due to F3, $b_{j}$ is placed close to the root, allowing $x_{j}$ to be almost any leaf. Consequently, $x_{i}$ remains independent of $x_{j}$.
@@ 29 $\mathcal{S}^{*}\leftarrow(\mathcal{S}^{c}\setminus\mathcal{S})\cup\{S\}$
The Strong MVP-ORAM protocol makes access pattern oblivious even in the worst-case scenario when ORAM accesses follow a uniform distribution by executing $\sigma\geq 0$ dummy accesses for each real access. The proposed approach creates a scheduler of accesses containing $\sigma$ dummy accesses and a single real access, such that in each timestep, no two clients will access the same block if $c\leq\sigma+1$. To construct this scheduler, each client specifies the address it will access when invoking $\mathsf{getPM}$, and they utilize this information to create the scheduler. Note that the address is encrypted; however, for simplicity, we omit the encryption details from the algorithm specification. The modified client and server specifications are presented in Algorithm 4 and Algorithm 5, respectively.
The implementation of this strategy requires changing a few lines in the MPV-ORAM protocol. The first change is the addition of the list $\mathcal{A}$ in the server. For each client $c_{i}$, the server stores in $\mathcal{A}$ the encrypted address that $c_{i}$ will access and a counter that tracks the number of accesses executed by $c_{i}$ (A5, L7-10). The address is used by $c_{i}$ to determine the timestep at which it will perform the real access, while the counter is used by the server to remove $c_{i}$’s access context on the $(\sigma+1)^{\mathit{th}}$ access’s eviction (A5, L18-20).
@@
Using $\tau_{\mathit{real}}$, $c_{i}$ verifies twice whether the current timestep $\tau$ matches $\tau_{\mathit{real}}$ during the access. If the check returns true, $c_{i}$ executes the real access by selecting a random path that passes through the slot where the block is stored (A4, L9-10) and applies the operation of $\mathit{read}$ or $\mathit{write}$ on the block (A4, L15-20). Otherwise, it selects a random path (A4, L12) to execute dummy access. Note that if the number of concurrent clients $c$ accessing the same address is more than $\sigma+1$, then $c-\sigma$ clients will have to execute real access in the same last timestep. Hence, if $c\leq\sigma+1$, then this strategy guarantees obliviousness even when the worst-case access pattern, i.e., all clients are concurrently accessing the same address, is not skewed. Next, we prove this result.
@@ E-A Obliviousness of Strong MVP-ORAM
Given $c,\sigma,N\in\mathbb{N}$, $\alpha\in\mathbb{R}$, and $\mathcal{D}\in\mathcal{U}$, the statistical distance between a random sequence of size $c$ and the access pattern generated by Strong MVP-ORAM under the dummy access strategy is bounded by $\mu(N,c,\mathcal{D}(\alpha),\sigma)$.
@@
  • If $c\leq\sigma+1$. Since clients can detect whether $\mathit{addr}$ is being accessed by others during timestep $t$, a client that is the $k$th to access $\mathit{addr}$, where $k\leq c$, performs k1k-1 dummy accesses before issuing its actual access. Indeed, the mitigation strategy ensures that accesses to $\mathit{addr}$ are performed so that no two clients access $\mathit{addr}$ in the same timestep. As each block is moved to the stash at the end of every access (if it is not already there), each client selects a random, independent path when accessing the block. As a result, the access pattern generated by Strong MVP-ORAM is indistinguishable from a random sequence of size $c$.
@@
Given $c,\sigma,N\in\mathbb{N}$, $\alpha\in\mathbb{R}$, and $\mathcal{D}\in\mathcal{U}$, let $\overrightarrow{b}_{e}=\{b_{i}\}_{i\in\{1,\dots,c\}}$ denote a set of $c$ concurrent queries in timestep $e$ and $\overrightarrow{y}=(\overrightarrow{b}_{1},\overrightarrow{b}_{2},\dots)$ denote a set of queries from multiple clients in each timestep. MVP-ORAM holds the Obliviousness property of a Wait-Free ORAM, as per Definition 1, when clients perform $\sigma$ dummy accesses.
@@
Let $A(\overrightarrow{y})$ be an access pattern generated by Strong MVP-ORAM while executing query $\overrightarrow{y}$. According to Definition 1, we must show that there is a function $\mu$ that bounds the statistical distance between any two access patterns of the same length. To establish such a bound, we compare $A(\overrightarrow{y})$ with a random access pattern. From Lemma 7, $A(\overrightarrow{y})$ is indistinguishable from a random access pattern of the same size when a single client accesses the ORAM in each timestep or multiple clients access different blocks. From Lemma 12, in the worst-case scenario, the statistical distance between $A(\overrightarrow{b}_{e})$ and a random access pattern of the same size is bounded by $\mu(N,c,\mathcal{D}(\alpha),\sigma)$. Since after each concurrent access ORAM changes, i.e., locations of blocks accessed by clients change, the statistical distance between $A(\overrightarrow{y})$ and a random access sequence is still bounded by $\mu(N,c,\mathcal{D}(\alpha),\sigma)$, proving the lemma. ∎
To conclude our obliviousness analysis, we show under which conditions MVP-ORAM can provide the same security as a traditional parallel ORAM with collision-freedom [boyle2015].
@@
When there are at most $\sigma+1$ clients accessing ORAM, they retrieve blocks in $\sigma+1$ timesteps. When they are retrieving different blocks, they will access different random paths in the first timestep and continue accessing random paths in the remaining $\sigma$ timesteps. When they retrieve the same block, each client will perform the real access in different timesteps and dummy accesses in the other ones. In both cases, clients generate a sequence of leaves by selecting any of the $2^{L}$ leaves in each access. Thus, the probability of all clients accessing the same leaf is $2^{-L\,c\,(\sigma+1)}$, which is the same probability of clients selecting random leaves and is negligible in $N$. ∎
@@ Appendix F Artifact Appendix
The artifact includes the implementation of MVP-ORAM, QuORAM, [^4] https://github.com/SeifIbrahim/QuORAM/ and COBRA. [^5] https://github.com/bft-smart/cobra The former was used for performance assessment, while the latter two were used for comparison. The experiments were conducted on multiple AWS servers, but for ease of reproduction, we provide a scaled-down version of a single-machine evaluation setup in this appendix.
@@ F-A1 How to access
The artifact is available on Zenodo, [^6] https://doi.org/10.5281/zenodo.17842154 which includes the MVP-ORAM, QuORAM, and COBRA implementations adapted for easier benchmarking.
@@ F-A3 Software dependencies
Execution of this artifact requires unzip, gnuplot, and OpenJDK 11. It can be executed on either Linux or the Linux subsystem on Windows. [^7] https://learn.microsoft.com/windows/wsl
@@ F-D Major Claims
  • (C1): The stash size stabilizes after some number of accesses, and the maximum stash size increases as we decrease $\alpha$ and increase $c$. This is proven by experiment (E1), whose results are shown in Fig. 7.
@@
  • (C4): MVP-ORAM outperforms QuORAM both in throughput and latency with $50$ clients. This is proven by experiment (E3), with results reported in Table IV.
  • (C5): MVP-ORAM’s performance is an order of magnitude lower than COBRA with $50$ clients. This is proven by experiment (E4), with results reported in Table IV.
  • (C6): The statistical distance between access patterns decreases as the parameter $\alpha$ and $L$ increase, i.e., accesses are more skewed and tree size increases. This is demonstrated by experiment (E5), whose results are shown in Fig. 5.
  • (C7): The statistical distance between access patterns decreases as the parameter $\sigma$ increases, confirming that MVP-ORAM’s security improves with higher $\sigma$ values. This is demonstrated by experiment (E6), whose results are shown in Fig. 6.
@@ F-E1 Experiment (E1)
  • controller.benchmark.strategy=oram. benchmark.MeasurementBenchmarkStrategy
@@
[Execution] Consider the project MVP-ORAM and follow the steps described in the experimental workflow to execute the experiment twice, setting the parameters defined above.
@@ F-E2 Experiment (E2)
  • controller.benchmark.strategy=oram. benchmark.MeasurementBenchmarkStrategy
@@
[Execution] Consider the project MVP-ORAM and follow the steps described in the experimental workflow to execute the experiment using the above values.